# Slice-Level Scheduling for High Throughput and Load Balanced LLM Serving

### v1

@@ 3.3. Actual Generation Length Distribution
Figure 7. SCLS system overview. The arrows represent data transferred between modules, where ❶ is requests, ❷ is temporary batches, ❸ and ❼ are estimated serving time, ❹ is estimated memory consumption, ❺, ❻, and ❽ are batches, ❾ is uncompleted requests, ❿ is completed requests.
4. Solution Description
4.1. System Overview
@@
SCLS has four core modules, a serving time estimator, a memory usage estimator, an adaptive batcher, and a max-min offloader. The system overview of SCLS is presented in Fig. 7.
@@
SCLS periodically fetches all requests from the request pool with a time interval $T$ and hands them over to the adaptive batcher. The adaptive batcher leverages a dynamic programming-based algorithm to group the requests into batches with the goal of minimizing the total serving time of all batches. In each step of the batching algorithm, the adaptive batcher leverages the estimated serving time of temporarily grouped batches to update the state transfer function and utilizes estimated memory consumption to guarantee that the serving process will not cause OOM errors. After the batching procedure finishes, batches will be offloaded to workers by the max-min offloader, who schedules batches with long estimated serving time to lightly loaded workers to achieve load balancing. After workers finish batch serving, completed requests are returned to the user, and uncompleted requests are sent to the request pool to be scheduled with newly arrived requests.
@@
A worker is an LLM instance running on a specific inference engine. The worker’s receiving thread receives the batch from the offloader and saves it in a local queue, while the worker’s processing thread continuously fetches a batch from the queue and serves it. In this paper, we implement the hf-worker and ds-worker to load and run LLM instances using hugging-face transformers and deepspeed-inference as inference engines, respectively.
4.2. Efficient Serving Time Estimation
The serving time estimator estimates the batch serving time according to the batch input length, batch size, and slice length.
Existing inference engines such as huggingface-transformers and deepspeed-inference serve requests with static batching. For each batch, they pad all the batched requests to the longest input length of requests in the batch before serving. Besides, during batch serving, if a request has generated the EOS token but there are other requests that need to continue generating tokens, it will still participate in the subsequent computation and generate invalid tokens which are discarded after batch serving. Therefore, for static batching, the serving time of a batch is determined by the batch size, the batch input length, and the batch generation length.
An intuitive way to estimate the batch serving time is to directly profile the serving time for various batch input lengths, batch generation lengths, and batch sizes. However, this is computationally expensive and requires additional profiling when the range of the predefined maximal generation length limit changes. To avoid this problem, we look at the serving procedure and model the time consumed in the prefill and decoding phases, respectively.
@@ 4.2. Efficient Serving Time Estimation
Figure 8. Prefill latency under deepspeed-inference (DS).
(a) $\tau_{decode}(l,N)$ under different cached lengths.
@@
Figure 9. Per-iteration decoding latency under DS.
@@
(1) $T_{serve}(N,L_{i},L_{o})=T_{prefill}(N,L_{i})+T_{decode}(N,L_{i},L_{o}),$
where $T_{prefill}(N,L_{i})$, and $T_{decode}(N,L_{i},L_{o})$ are the serving latency in the prefill and decoding phase, respectively. Since the prefill phase computes key and value tensors for all the input tokens, the latency is determined by the batch input length $L_{i}$ and the batch size $N$. In the decoding phase, the serving procedure runs in an iterative manner, and therefore $T_{decode}(N,L_{i},L_{o})$ can be calculated by summing up the latency in each decoding iteration, which is expressed by
(2) $T_{decode}(N,L_{i},L_{o})=\sum_{l=1}^{L_{o}}\tau_{decode}(L_{i}+l,N),$
where $\tau_{decode}(l,N)$ denotes the serving latency in the decoding iteration. Since in each decoding iteration, for each request, only one token is fed to the LLM and extracts features from all the previously cached tokens to make a prediction, $\tau_{decode}(l,N)$ is determined by the cached length $l$ and the batch size $N$.
We profile the prefill latency $T_{prefill}(N,L_{i})$ of an LLaMA2-13B instance using deepspeed-inference-v0.13.3 as the inference engine under various batch sizes and batch input lengths. To clearly show the individual effect of batch size $N$ and batch input length $L_{i}$ on $T_{prefill}(N,L_{i})$, we plot two sub-figures with these two factors as the horizontal axis in Fig. 8. Similarly, we also profile the per-iteration latency $\tau_{decode}(l,N)$ for different cached lengths and batch sizes and show the effect of cached length $l$ and batch size $N$ on $\tau_{decode}(l,N)$ in Fig. 9(a) and Fig. 9(b), respectively.
@@
Figure 10. Serving time estimation error of various inference engines, where HF and DS are short for huggingface-transformers and deepspeed-inference, respectively.
@@
For the prefill latency $T_{prefill}(N,L_{i})$, we can observe that when the batch size is fixed, $T_{prefill}(N,L_{i})$ increases linearly with the increase of the batch input length $L_{i}$ from Fig. 8(a). Besides, as shown in Fig. 8(b), when the input length is fixed, the $T_{prefill}(N,L_{i})$ increases linearly with the increase of the batch size $N$ as well. We also profile the prefill latency for huggingface-transformers, and observe the same phenomenon. Therefore, we use the following function to fit $T_{prefill}(N,L_{i})$.
(3) $T_{prefill}(N,L_{i})=p_{1}\cdot N\cdot L_{i}+p_{2}\cdot N+p_{3}\cdot L_{i}+p_{4},$
@@
where $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are parameters to be fitted. Under this function, $T_{prefill}(N,L_{i})$ is a linear function of the batch input length $L_{i}$ when the batch size $N$ is fixed, and when $L_{i}$ is fixed, $T_{prefill}(N,L_{i})$ is also a linear function of $N$, which is consistent with our observations. We leverage the profiled prefill latency data of various engines to fit the function. Fig. 10(a) presents the fitting error metric, root-mean-square error (RMSE) which represents the average gap between the estimated and actual latency. As shown in Fig. 10(a), the average gap is merely 0.16s for huggingface-transformers and less than 0.04s for deepspeed-inference, achieving an accurate estimation. Hence, Eq. (3) is used to compute the prefill latency.
@@
For the per-iteration decoding latency $\tau_{decode}(l,N)$. we can observe from Fig. 9(a) that when the batch size is fixed, $\tau_{decode}(l,N)$ rises with fluctuations as the cached length increases. Besides, as shown in Fig. 9(b), when the cached length is fixed, $\tau_{decode}(l,N)$ also shows a rising trend with the increase of the batch size, and when the cached length is large, the tendency tends to be linear. We also collect per-iteration decoding latency profiles for huggingface-transformers, and observe the same phenomenon. We try to estimate the per-iteration latency the same way we estimate the prefill latency using the following function.
(4) $\tau_{decode}(l,N)=d_{1}\cdot N\cdot l+d_{2}\cdot N+d_{3}\cdot l+d_{4},$
@@
where $d_{1}$, $d_{2}$, $d_{3}$, and $d_{4}$ are parameters to be fitted. Fig. 10(a) presents the estimation error of $\tau_{decode}(l,N)$ under various inference engines, we can find that the fitting error is negligible, which confirms that the estimation is accurate.
@@
Thus, by incorporating Eq. (4) into Eq. (2), we can obtain the estimated serving time of the decoding phase. Adding the decoding latency to the prefill latency computed by Eq. (3), we can get an estimate of the batch serving time. However, although the estimation error of a single decoding iteration is low, the error accumulates as the generation process goes on. Therefore, we evaluate the estimation error of overall batch serving time for an iteration number of 128 under various batch input lengths and batch sizes. As shown in Fig. 10(b), the average error is merely 2.3s for HF and less than 0.4s for DS, which confirms that the estimation is accurate.
@@
From Fig. 10, we can find that both the per-iteration error and multi-iteration error of HF are much higher than those of DS. This is because DS leverages customized CUDA kernels to accelerate the computation of transformer blocks, so its latency bases are much smaller than that of HF, and hence the average estimated latency error is smaller than that of HF.
@@
SCLS sets the iteration number limit for static batching to a small slice length $S$, when the number of iterations reaches $S$, the batch serving ends regardless of whether or not all requests have generated the EOS token. Therefore, we can consider the batch generation length $L_{o}$ to be $S$ for batch serving. Hence, the serving time of SCLS can be accurately and quickly estimated by $T_{serve}(N,L_{i},S)$, where $N$ and $L_{i}$ can be directly obtained before serving.
Compared with estimating the batch serving time by directly profiling the whole serving time under various batch input lengths and batch sizes for a given slice length, such a modeling approach adopted by the serving time estimator only needs to profile the prefill and decoding latency of a single iteration, which greatly reduces the computational overhead of profiling. Moreover, when the slice length changes, the serving time estimator does not need re-profiling, which achieves a flexible and efficient estimation.
@@
If all requests in the batch generate the EOS token before the generation length reaches $S$, the batch serving will finish prematurely and the estimated serving time will be inaccurate. However, since $S$ is always set small, the likelihood of the premature finish happening is small, and the performance of SLCS won’t be affected. We also verify this through experiments in Section 5.
4.3. Practical Memory Usage Estimation
The memory usage estimator leverages the batch size, batch input length, and slice length to determine whether a batch of requests will cause OOM errors during batch serving.
@@ 4.3. Practical Memory Usage Estimation
During LLM serving, GPU memory is mainly occupied by the parameters of the LLM and the key-value cache of requests. The memory consumed by the LLM parameters is constant, but the memory consumed by the key-value cache varies with the batch size, batch input length, and batch generation length. Since for static batching, both the pad tokens and the invalid tokens generated in the request-waiting process produce key and value tensors that are cached and occupy memory, the memory consumption of the key-value cache can be calculated by
(5) $M_{kv}(N,L_{i},L_{o})=(L_{i}+L_{o})\cdot N\cdot\Delta,$
@@
where $N$, $L_{i}$, $L_{o}$, and $\Delta$ respectively represent the batch size, the batch input length, the batch generation length, and the per-token memory usage of key and value tensors. $\Delta$ is determined by the model architecture.
@@
(6) $M_{ava}=M_{cap}-M_{model}-M_{engine},$
where $M_{cap}$ is the GPU memory capacity, $M_{model}$ is the memory occupied by the LLM parameters, and $M_{engine}$ is the memory utilized by the inference engine to store engine-specific data.
@@
(7) $M_{kv}(N,L_{i},S)\leq M_{ava}.$
In addition, by incorporating Eq. (5) into Eq. (7) and rearranging it, we can obtain Eq. (8), which shows that for a given input length $L_{i}$, the maximal batch size $N_{max}$ can be smaller when the slice length is longer. Therefore, if we set the slice length $S$ to the predefined maximal generation length, $N_{max}$ will be small and SLCS degenerates into SLS. On the contrary, by setting the slice length $S$ to a small value, more requests can be batched together to achieve a high throughput.
(8) $N_{max}=\lfloor\frac{M_{ava}}{\Delta\cdot(L_{i}+S)}\rfloor.$
In practice, for the same model and batched requests, there is a gap between the actual GPU memory usage and the theoretical memory usage for different inference engines due to the engine-specific memory management mechanism.
Huggingface-transformers (HF), as an inference engine based entirely on pytorch (paszke2019pytorch, ), produces memory fragmentation that cannot be utilized during the batch serving. Therefore, for HF, it is necessary to multiply $M_{ava}$ in Eq. (7) by an additional coefficient $\zeta$ less than 1, and the constraint is rewritten as
(9) $M_{kv}(N,L_{i},S)\leq\zeta\cdot M_{ava}.$
Deepspeed-inference (DS) has an inflexible memory management mechanism. Although it can serve batched requests with a larger batch size when the batch input and generation lengths are shorter, we experimentally find that it is unable to determine whether OOM errors will occur using Eq. (9), even if $\zeta$ is set to a small value such as 0.6. Therefore, for DS, under a given slice length $S$, we profile the maximal batch size for various batch input lengths and make rules for determining whether OOM will occur during batch serving.
4.4. Serving Time-Oriented Batching
@@
Figure 11. Batching example. Under a given slice length of 128, various batching strategies have a large difference in the serving time. In sub-figure (a), white and blue grids represent pad and input tokens, respectively.
As shown in Fig. 11, under a slice length of 128, given 15 requests with an input length of 10 and 1 request with an input length of 1024, if they are batched together, it takes 13.5s to finish serving for an LLaMA2-13B instance deployed with huggingface-transformers-v4.35.0. If the requests with input lengths of 10 and 1024 are divided into two separate batches, the total serving time is only 7.6s, which is far less than the serving time of batching them together. This is because all the requests are padded to 1024 tokens when they are batched together, which greatly increases the per-iteration prefill and decoding latency for the requests with an input length of 10.

### v2

@@ III-B Unbalanced Workload of SLS and ILS
Figure 4: SCLS system overview. The arrows represent data transferred between modules, where ❶ is requests, ❷ is temporary batches, ❸ and ❼ are estimated serving time, ❹ is estimated memory consumption, ❺, ❻, and ❽ are batches, ❾ is uncompleted requests, and ❿ is completed requests.
III-C Real-World Generation Length Distribution
To understand the generation length distribution in real-world LLM-based web applications, we analyze 400,000 pieces of data from ShareGPT [^1] ShareGPT [11] is a publicly available website that shares conversations from users chatting with ChatGPT [8]. and one-month request traces collected from CodeFuse’s logs. We depict the distribution of request input and generation lengths in Fig. 3, where we can find that the vast majority of requests have a small generation length of less than 512 in both CodeFuse and ShareGPT data. Hence, if we split the maximal generation length limit (e.g., 2048) into slices with a small slice length (e.g., 128) and serve batched requests slice by slice, most requests can be returned in time within a small number of slices. In addition, since the iteration number limit is set to the small slice length, we can get a precise range of memory consumption and batch together as many requests as possible while ensuring that OOM errors will not occur, thus increasing throughput.
Moreover, for static batching, we can also get a precise range of batch serving time. Hence, load balancing can be achieved by offloading batches with the max-min policy where the batch with the longest serving time is offloaded to the least loaded instance.
Considering the above, we propose SCLS, which splits the predefined maximal generation length limit into fixed-length slices, limits the number of iterations in each batch serving to the slice length, and serves requests slice by slice. Although requests with long generation lengths suffer from the extra overhead of recomputing the prefill phase at every reschedule, the overhead is acceptable because these long-generated requests are rare as shown in Fig. 3.
IV Solution Description
IV-A System Overview
@@
SCLS has four core components, a serving time estimator, a memory usage estimator, an adaptive batcher, and a max-min offloader. The overview of SCLS is presented in Fig. 4.
@@
SCLS periodically fetches all requests from the request pool with a time interval $T$ and hands them over to the adaptive batcher. The adaptive batcher leverages a dynamic programming-based algorithm to group the requests into batches with the goal of minimizing the total serving time. In each step of the batching algorithm, the adaptive batcher leverages the estimated serving time of temporarily grouped batches to update the state transfer function and utilizes estimated memory consumption to guarantee that the serving process will not cause OOM errors. After the batching procedure finishes, batches will be offloaded to workers by the max-min offloader. After batch serving is finished, completed requests are returned to the user, and uncompleted requests are sent to the request pool to be scheduled with newly arrived requests.
@@
A worker in Fig. 4 is an LLM instance running on a specific inference engine. The worker’s receiving thread receives the batch from the offloader and saves it in a local queue, while its processing thread continuously fetches a batch from the queue and serves it. In this paper, we implement the HF-worker and DS-worker to load and run LLM instances using huggingface transformers (HF) [12] and deepspeed-inference (DS) [9] as inference engines, respectively.
@@
Figure 5: Prefill latency under deepspeed-inference (DS).
(a) $\tau_{decode}(l,N)$ under different context lengths.
@@
Figure 6: Per-iteration decoding latency under DS.
IV-B Efficient Serving Time Estimation
Inference engines, such as HF and DS, serve requests with static batching, where the batch serving time is determined by the batch size, batch input length, and batch generation length.
@@ IV-B Efficient Serving Time Estimation
$T_{serve}(N,L_{i},L_{o})=T_{prefill}(N,L_{i})+T_{decode}(N,L_{i},L_{o}),$ (1)
where $T_{prefill}(N,L_{i})$, and $T_{decode}(N,L_{i},L_{o})$ are the serving latency in the prefill and decoding phase, respectively. Since the prefill phase computes key and value tensors for all the input tokens, the latency is determined by the batch input length $L_{i}$ and the batch size $N$.
In the decoding phase, the serving procedure runs iteratively. Therefore, $T_{decode}(N,L_{i},L_{o})$ can be calculated by summing up the per-iteration decoding latency, which is expressed by
$T_{decode}(N,L_{i},L_{o})=\sum_{l=L_{i}+1}^{Li+L_{o}}\tau_{decode}(l,N),$ (2)
where $\tau_{decode}(l,N)$ denotes the per-iteration latency. In each decoding iteration, for each request, only one token is fed to the LLM and extracts features from all the previously cached tokens to make a prediction, and hence $\tau_{decode}(l,N)$ is determined by the context length $l$ and the batch size $N$.
We profile the prefill latency $T_{prefill}(N,L_{i})$ of an LLaMA2-13B instance using DS as the inference engine under various batch sizes and batch input lengths. To clearly show the individual effect of batch size $N$ and batch input length $L_{i}$ on $T_{prefill}(N,L_{i})$, we plot two sub-figures with these two factors as the horizontal axis in Fig. 5. Similarly, we also profile the per-iteration latency $\tau_{decode}(l,N)$ and show the effect of context length $l$ and batch size $N$ in Fig. 6.
@@
Figure 7: Serving time estimation error of various inference engines, where HF and DS are short for huggingface-transformers and deepspeed-inference, respectively.
@@
For the prefill latency $T_{prefill}(N,L_{i})$, we can observe that when the batch size is fixed, $T_{prefill}(N,L_{i})$ increases linearly with the batch input length $L_{i}$ from Fig. 5(a). Besides, as shown in Fig. 5(b), when the input length is fixed, the $T_{prefill}(N,L_{i})$ increases linearly with the batch size $N$ as well. For HF, we observe the same phenomenon. Therefore, we use the following function to fit $T_{prefill}(N,L_{i})$.
$T_{prefill}(N,L_{i})=p_{1}\cdot N\cdot L_{i}+p_{2}\cdot N+p_{3}\cdot L_{i}+p_{4},$ (3)
@@
where $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are parameters to be fitted. Under this function, $T_{prefill}(N,L_{i})$ is a linear function of the batch input length $L_{i}$ when the batch size $N$ is fixed. Besides, when $L_{i}$ is fixed, $T_{prefill}(N,L_{i})$ is also a linear function of $N$, which is consistent with our observations. We leverage the profiled prefill latency data of various engines to fit the function. Fig. 7(a) depicts the fitting error metric, root-mean-square error (RMSE), which represents the average gap between the estimated and actual latency. As shown in Fig. 7(a), the average gap is merely 0.16s for HF and less than 0.04s for DS, achieving an accurate estimation.
@@
For the per-iteration decoding latency $\tau_{decode}(l,N)$. we can observe from Fig. 6(a) that when the batch size is fixed, $\tau_{decode}(l,N)$ rises with fluctuations as the context length increases. Besides, as shown in Fig. 6(b), when the context length is fixed, $\tau_{decode}(l,N)$ also shows a rising trend with the increase of the batch size, and when the context length is large, the tendency tends to be linear. For HF, we also observe the same phenomenon. Hence, we try to estimate the per-iteration latency the same way we estimate the prefill latency with
$\tau_{decode}(l,N)=d_{1}\cdot N\cdot l+d_{2}\cdot N+d_{3}\cdot l+d_{4},$ (4)
@@
where $d_{1}$, $d_{2}$, $d_{3}$, and $d_{4}$ are parameters to be fitted. Fig. 7(a) presents the estimation error of $\tau_{decode}(l,N)$ for various inference engines. We can find that the fitting error is negligible, which confirms that the estimation is accurate.
@@
By incorporating Eq. (4) into Eq. (2), we can obtain the estimated serving time of the decoding phase. Adding the decoding latency to the prefill latency computed by Eq. (3), we can get an estimate of the batch serving time. We also evaluate the estimation error of overall batch serving time for an iteration number of 128 under various batch input lengths and batch sizes. As shown in Fig. 7(b), the average error is merely 2.3s for HF and less than 0.4s for DS, which confirms that the estimation is accurate.
@@
From Fig. 7, we can find that both the per-iteration error and multi-iteration error of HF are much higher than those of DS. This is because DS leverages customized CUDA kernels to accelerate the computation of transformer blocks, so its latency bases are much smaller than that of HF, and hence its average estimation error is smaller than that of HF. Moreover, since the prefill latency is typically much larger than the per-iteration decoding latency, the average estimation error of the prefill latency is also larger than that of the decoding latency.
@@
SCLS sets the iteration number limit for static batching to a small slice length $S$. When the number of iterations reaches $S$, the batch serving ends regardless of whether or not all requests have generated the EOS token. Therefore, we can consider the batch generation length $L_{o}$ to be $S$ for batch serving. Hence, the batch serving time of SCLS can be accurately and quickly estimated by $T_{serve}(N,L_{i},S)$, where $N$ and $L_{i}$ can be directly obtained before batch serving.
@@
If all requests in the batch generate the EOS token before the generation length reaches $S$, the batch serving will finish prematurely and the estimated serving time will be inaccurate. However, since $S$ is always set small, the likelihood of the premature finish happening is small, and the performance of SLCS won’t be affected. We verify this in Section V.
The serving time estimator is trained on offline profiled data and then applied for online prediction. Compared with exhaustively profiling the serving time under various batch input lengths and batch sizes for a given slice length, such an analytical model can be established merely by profiling the per-iteration prefill and decoding latency, greatly reducing the profiling overhead. Furthermore, the online estimation is also simple and efficient.
IV-C Practical Memory Usage Estimation
@@
During LLM serving, GPU memory is mainly occupied by the parameters of the LLM and the key-value cache of requests. The memory consumed by the key-value cache is determined by the batch size, batch input length, and batch generation length. Since for static batching, both the pad tokens and generated invalid tokens produce key-value cache, the memory consumed by key-value cache is calculated by
$M_{kv}(N,L_{i},L_{o})=(L_{i}+L_{o})\cdot N\cdot\Delta,$ (5)
@@
where $N$, $L_{i}$, and $L_{o}$ respectively represent the batch size, the batch input length, and the batch generation length. $\Delta$ denotes the per-token memory usage of key and value tensors, which is determined by the model architecture.
@@
$M_{ava}=M_{cap}-M_{model}-M_{engine},$ (6)
where $M_{cap}$ is the GPU memory capacity, $M_{model}$ is the memory consumed by LLM parameters, and $M_{engine}$ is the memory utilized by the inference engine to store engine-specific data.
@@
$M_{kv}(N,L_{i},S)\leq M_{ava}.$ (7)
In addition, by incorporating Eq. (5) into Eq. (7) and rearranging it, we can obtain Eq. (8), which indicates that given the batch input length $L_{i}$ and slice length $S$, limiting the batch size to be less than $N_{max}$ can avoid OOM errors.
$N_{max}=\lfloor\frac{M_{ava}}{\Delta\cdot(L_{i}+S)}\rfloor.$ (8)
In practice, there is a gap between the actual GPU memory usage and the theoretical usage for different inference engines due to the engine-specific memory management mechanism.
HF is an inference engine based entirely on pytorch [13]. It produces memory fragmentation that cannot be utilized during the batch serving. Therefore, for HF, it is necessary to multiply $M_{ava}$ in Eq. (7) by an additional coefficient $\zeta$ less than 1, and the constraint is rewritten as
$M_{kv}(N,L_{i},S)\leq\zeta\cdot M_{ava}.$ (9)
DS has an inflexible memory management mechanism. We experimentally find that it is unable to determine whether OOM errors will occur using Eq. (9) for DS, even if $\zeta$ is set to a small value. Therefore, for DS, for a given slice length $S$, we profile the maximal batch size for various batch input lengths and make rules for determining whether OOM will occur during batch serving.
IV-D Serving Time-Oriented Batching
@@
Figure 8: Batching example. Under a given slice length of 128, various batching strategies have a large difference in the serving time. In sub-figure (a), white and blue grids represent pad and input tokens, respectively.
As shown in Fig. 8, under a slice length of 128, given 15 requests with an input length of 10 and 1 request with an input length of 1024, if they are batched together, it takes 13.5s to finish serving for an LLaMA2-13B HF-worker deployed on one NVIDIA A100 80GB GPU. If the requests with input lengths of 10 and 1024 are divided into two separate batches, the total serving time is only 7.6s. This is because all the requests are padded to 1024 tokens when they are batched together, which greatly increases the per-iteration prefill and decoding latency for the requests with an input length of 10.
@@
Pseudocode of the batching algorithm is presented in Algorithm 1. It first sorts requests in ascending order by their input lengths. Then the state array $\mathcal{T}$ is created, where $\mathcal{T}[i]$ denotes the minimal total batch serving time for the first $i$ requests. The initial state $\mathcal{T}[0]=0$ indicates that the total serving time is 0 when there is no request to batch. The position array $\mathcal{P}$ stores positions for partitioning requests into batches.
During dynamic programming, the algorithm sequentially traverses each request. For each request, the algorithm first treats it as a separate batch and then tries to batch it together with preceding requests to minimize the total batch serving time while guaranteeing that OOM errors will not occur. The state transfer function of dynamic programming is
$\mathcal{T}[i]=\min\limits_{0<j\leq i}(\mathcal{T}[j-1]+T_{serve}(i-j+1,L_{i},S)),$ (10)
where $i-j+1$ computes the batch size. Since requests are sorted by the request input length at the beginning when traversing to the $i$th request, its input length is the maximal input length of the first $i$ requests, and hence its input length can be utilized as the batch input length $L_{i}$ to calculate the estimated batch serving time $T_{serve}$. After the dynamic programming, the requests are divided into batches according to the recorded splitting positions.
In the batching procedure, whether OOM errors will occur is determined by the memory usage estimator, and the batch serving time is estimated by the serving time estimator.